# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v10

@@ 4.3.2 Chow-Liu Approximation
To handle these scaling issues, we adopt Chow–Liu trees chow1968approximating for approximating a high-dimensional joint distribution using a maximum-weight spanning tree, $\pi$, over the different sensor modalities. Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies. This approach minimises the Kullback–Leibler divergence (Def. 2.2) between the true multivariate distribution and the resulting tree-based approximation as follows:
@@ 4.3.3 Results
Tables 47 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low. For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes barely reduces min-entropy at all. For the Relay dataset in Table 5, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the $8.092$ bits produced by the full of set of sensors. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
@@ 4.3.4 Binning Sensitivity Analysis
As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently. Too few bins cause oversmoothing, erasing distributional features and deflating entropy. At the same time, too many bins induces sparsity and overfitting, inflating entropy by modelling noise rather than the source. We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets. We re-calculated the joint entropy while varying the number of bins for each modality, starting from 5 bins and increasing up to 2048. Beyond that, the computational cost of processing the large state space became prohibitive. The results of this analysis are presented in Figure 4.
The graphs show a consistent trend across all datasets. For a low number of bins, all entropy measures are severely underestimated as expected. As the number of bins increases, the entropy rises sharply before tapering off and converging towards a stable plateau that approximates a logarithmic shape. (An exception is the min-entropy value for SHL, which converges after approximately 100 bins). Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
@@
Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) values are also shown.
@@ 5 Evaluation
This section interprets our entropy findings for system design: we discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
@@ 5.1 Discussion
Single sensors yield low entropy. Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g. some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
Combining multiple sensors offers modest gains with respect to worst-case min-entropy. While using multi-modal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect. In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes. Even the most complex multi-modal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards. This also aligns with prior work showing that Shannon entropy can significantly overstate security arikan2002inequality , cachin1997entropy , turan2018recommendation , bsi2024ais31 . Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves. Using multiple sensors may still improve usability and model accuracy, e.g. for coarse co-location checks, but they should not be treated as secure, unpredictable sources.
@@ 5.2 Translating Entropy Bounds Into Guessing Attacks
We now sketch how low-entropy sensor data translates into concrete attack cost. Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy. Ordering guesses by decreasing probability, the success probability after $q$ guesses satisfies
@@
The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ cachin1997entropy .
Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching mehrnezhad2015tap , gurulian2018good , shrestha2014drone , markantonakis2024using , alzubaidi2016authentication , mayrhofer2021adversary , abuhamad2020sensor . Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$. An adversary’s task becomes easier through a similarity-based scheme: they only need to generate a candidate $z^{\prime}$ that is ‘close enough’ to a high-probability vector $z$. This means the effective success probability for each guess is, in fact, higher than the bound in Eq. 10 suggests. The best-performing combination in our study yielded $H_{\infty}\approx 24$ bits, corresponding to a search space of $2^{24}$ ($\approx$16.7 million) possibilities. An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average. This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys nistsp80057 . Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach mai2017guessability . In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average. A main defence against this is to employ rate-limit the number of attempts, although it does not solve the root problem. If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$. That is, rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
@@ 5.3 What can be done?
User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits voris2011accelerometers , lv2020analysis . For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
@@
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see suciu2011unpredictable , wallace2016toward ). Cryptographic extractors (e.g. Von Neumann and later constructions von195113 , trevisan2001extractors , barak2003true ) or decorrelation techniques (e.g. PCA) may be considered to convert low-quality sensor noise into higher-entropy sources. However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost. Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
@@ 5.4 Scope and Limitations
We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.). This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings. For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g. min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties). Thus, the unconditional entropy values we calculate in this work act as an upper bound. We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial. Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
@@ Acknowledgments
Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.

### v12

@@ 4.3.2 Chow-Liu Approximation
To handle these scaling issues, we adopt Chow–Liu trees chow1968approximating to approximate a high-dimensional joint distribution using a maximum-weight spanning tree, $\pi$, over the different sensor modalities. Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies. This approach minimises the Kullback–Leibler divergence (Def. 2.2) between the true multivariate distribution and the tree-based approximation as follows:
@@ 4.3.3 Results
Tables 58 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset. Combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy. However, min-entropy ($H_{\infty}$) remains stubbornly low; the complete set of sensors in SHL, for instance, surpasses 80 bits of $H_{1}$, but saturates at only 21 bits of $H_{\infty}$. Interestingly, we find that omitting certain correlated sensors sometimes barely reduces min-entropy at all. For the Relay dataset in Table 6, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the $8.092$ bits when using every sensor. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$. This pattern appears across datasets: additional modalities raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
@@ 4.3.4 Binning Sensitivity Analysis
As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently. Too few bins cause oversmoothing, erasing distributional features and deflating entropy. At the same time, too many bins induces sparsity and overfitting, inflating entropy by modelling noise rather than the source. We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets. We re-calculated the joint entropy while varying the number of bins for each modality, starting from 5 bins and increasing up to 2048. Beyond that, the computational cost of processing the large state space became prohibitive. The results of this analysis are presented in Figure 4. For completeness, we also give an alternative binning strategy—Scott’s rule ($3.5\sigma\cdot n^{-1/3}$)—as a comparator for our FD decision rule, which we report as a stability check.
The graphs show a consistent trend across all datasets. For a low number of bins, all entropy measures are severely underestimated as expected. As the number of bins increases, the entropy rises sharply before converging towards a stable plateau that approximates a logarithmic shape. (An exception is the min-entropy of SHL, which converges after $\approx$100 bins). Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
@@
Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) and Scott’s rule values are also shown.
@@ 5 Evaluation
We now discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
@@ 5.1 Discussion
Single sensors yield low entropy. Building on prior work, we find that sensor data has stubbornly low min-entropy (e.g. some single-sensor readings yield only 1–3 bits) and remains significantly below Shannon entropy. This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors—a critical distinction that existing schemes and evaluations do not capture at this scale.
Combining multiple sensors offers modest gains with respect to worst-case min-entropy. While using multi-modal sensors typically raises best- and average-case entropies, our experiments reveal that gains in worst-case min-entropy are far smaller than one might expect. In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes. Even the most complex multi-modal combinations in our datasets yield only modest worst-case entropy—between 8–24 bits in the best cases—which is well below the levels required by modern security standards. This also aligns with prior work showing that Shannon entropy can significantly overstate security arikan2002inequality , cachin1997entropy , turan2018recommendation , bsi2024ais31 . Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves. Using multiple sensors may still improve usability and model accuracy, e.g. for coarse co-location checks, but they should not be treated as secure, unpredictable sources.
@@ 5.2 Translating Entropy Bounds Into Guessing Attacks
We now sketch how low-entropy sensor data translates into concrete attack cost. Let $Z$ be a random variable over the discretised joint sensor measurement space $\mathcal{Z}$ used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy. If an adversary enumerates candidates $z\in\mathcal{Z}$ in order of decreasing probability $\Pr[Z=z]$, then after $q$ guesses the success probability is bounded by
@@
The expected number of guesses is $E[G]\approx 2^{H_{\infty}(Z)-1}$ cachin1997entropy in order to identify the correct sensor vector.
Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching mehrnezhad2015tap , gurulian2018good , shrestha2014drone , markantonakis2024using , alzubaidi2016authentication , mayrhofer2021adversary , abuhamad2020sensor . Let $\mathsf{Auth}:\mathcal{Z}\to\{0,1\}$ be an abstract authentication function that accepts ($1$) or rejects (0) an input vector $z\in\mathcal{Z}$. Under such a similarity-based scheme, the adversary does not need to reproduce the exact enrolled sensor vector; it is sufficient to generate some $z^{\prime}$ that lies within the acceptance region of a high-probability template. This means the effective success probability for each guess is higher than the bound in Eq. 10 suggests. The best-performing combination in our study yielded $H_{\infty}\approx 24$ bits, corresponding to a search space of $2^{24}$ ($\approx$16.7 million) possibilities. An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average. This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys nistsp80057 . Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach mai2017guessability . In the worst cases, such as single-sensor modalities where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring $\approx 2^{7}$ guesses on average. We shall now give more concrete indicative guesswork efforts in line with Cachin cachin1997entropy . Consider an adversary who has unlimited guess attempts against $\mathrm{Auth}$, which returns an immediate success/failure result. We assume the attacker submits $q$ queries, i.e. measurement(s), in descending order of their probability. The attacker effort can thus be characterised by the results given in Table 9. A primary defence here is to rate-limit the number of attempts; if each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$. Nevertheless, such limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
Table 9: Illustrating attacker effort versus min-entropy under optimal online guessing.
$H_{\infty}$ $E[G]$ (guesses) Time q=1q{=}1/s q=10q{=}10/s $q{=}10^{3}$/s $q{=}10^{6}$/s
8 $128$ $128$ s $12.8$ s $0.128$ s $0.128$ ms
12 2,0482{,}048 $34.1$ min $3.41$ min $2.05$ s $2.05$ ms
16 32,76832{,}768 $9.10$ h $54.6$ min $32.8$ s $3.28$ ms
20 524,288524{,}288 $6.07$ d $14.6$ h $8.74$ min $0.524$ s
24 $8.39\times 10^{6}$ $97.1$ d $9.71$ d $2.33$ h $8.39$ s
@@ 5.3 What can be done?
User actions such as prescribed shakes or gestures, or externally induced signals, can raise entropy by several bits voris2011accelerometers , lv2020analysis . For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors. A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities. Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
@@
Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs. The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see suciu2011unpredictable , wallace2016toward ). To improve the statistical quality of sensor-derived noise, one may apply cryptographic extractors (e.g. the Von Neumann extractor and newer schemes von195113 , trevisan2001extractors , barak2003true ) or decorrelation methods that attenuate dominant, deterministic trends. For example, principal component analysis (PCA) has been used elsewhere to isolate high-variance, predictable components—for example, gravitational bias in accelerometer axes—from low-variance residuals dominated by stochastic noise, which may then be discretised into candidate bitstrings (e.g. liu2025post ). The data processing inequality is relevant here, where $H_{\infty}(f(X))\leq H_{\infty}(X)$, for any deterministic transformation, $f$ vadhan2012pseudorandomness . That is, while extractors and transformations can reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more. Thus, such transformations are best interpreted as preconditioning or whitening steps that facilitate later modelling or extraction, but cannot increase the underlying min-entropy of the source itself. Future work should focus on designing sensor-based schemes that maximise shared entropy between devices while preserving the practical utility of sensor data for their intended applications.
@@ 5.4 Scope and Limitations
Throughout this work,, we report unconditioned, per-sample entropies for each sensor. This treats consecutive samples as independent and identically distributed (i.i.d.), which yields an upper bound on unpredictability because real sensor streams are time-correlated. For any additional side information, $Y$ (e.g. past samples), conditioning cannot increase entropy, i.e. $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties). We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial. Similarly, we may consider the entropy of a block of measurements over time. Under naïve i.i.d. conditions, and for intuition purposes only, the entropy rate of a source using block size $N$ and sampling frequency, $f_{s}$, can be considered at most $f_{s}\,\cdot\,N$; for the average modality in PerilZIS (3.912), this naïvely yields a maximum of $\approx$39 bits/sec under a 10Hz sampling frequency. Any temporal dependence can only reduce this rate (cf. §6, NIST SP 800-90B turan2018recommendation ). A full temporal analysis warrants a dedicated study, but we posit that its effects would be to decrease the figures reported in this work, reinforcing our conclusion that mobile sensors are weak sources of unpredictability for security applications.
@@ Acknowledgments
Carlton Shepherd received funding from the UK’s EPSRC (EP/Y030168/1) and would like to thank Darren Hurley-Smith for insightful conversations on the topic.
Appendix A Mutual Information Heatmaps
Figure A.1 shows the pairwise mutual information (in bits) between the sensors of each dataset.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure A.1: Pairwise mutual information (in bits) heatmaps for each dataset.